# Toricity in families of Fano varieties

### v1

@@ 1. Introduction
Let $\cal X\to T$ be a family of projective varieties over $\mathbb{C}$. The subset
$T_{\rm rational}\coloneqq\{\text{closed }t\in T\mid\mathcal{X}_{t}\text{ is a rational variety, i.e., $\mathcal{X}_{t}$ is birational to projective space}\}$
@@
It is expected that the previous theorem should hold without the Fano assumption. We note that being toric is neither an open nor a closed condition (Examples 6.1 and 6.2).
Being toric places strong restrictions on the geometry of a variety. For instance, smooth toric Fano varieties do not deform [2] (see also [11, Theorem 1.4]), so Theorem 1.1 is automatic in the smooth case. Thus, we consider an intermediate class: cluster type varieties (see Definition 2.10). This is a class of rational varieties generalizing toric varieties, and was introduced by Enwright, Figueroa, and the second author in [15]. For example, smooth toric varieties are covered by tori, whereas smooth cluster type Fano varieties contain an open dense torus (see [15, Theorem 1.2 and Theorem 1.3(4)]). We prove the following result about cluster type varieties in families of Fano varieties:
@@
By parametrizing, we mean here that every $n$-dimensional smooth cluster type Fano appears as the closed fiber of some $f_{i}$, and, moreover, every closed fiber of each $f_{i}$ is in this class of varieties. This notion is stronger than boundedness, but weaker than moduli, since the same variety could appear as a fiber multiple times.
Furthermore, from the proof of Theorem 1.2, it follows that all fibers of each $f_{i}$ in Corollary 1.3 admit the same combinatorial structure, in the sense that a certain birational transformation to a toric pair can be performed in the whole family (see Lemma 2.29).
For smooth rational Fano varieties, it is not clear whether the analogous statement to Corollary 1.3 should be true. In the toric case, the analogous statement was already known: smooth toric Fano varieties correspond to smooth reflexive polytopes (see, e.g., [9, Theorem 8.3.4]), so in each dimension $n$ there are only finitely many up to isomorphism. Note that smooth cluster type Fanos do deform, so unlike in the toric case, the parametrizing varieties $T_{i}$ in Corollary 1.3 are in general positive-dimensional. For example, every smooth del Pezzo surface of degree $\geq 2$ is of cluster type by [1, Theorem 2.1] and [22, Lemma 1.13].
@@ Acknowledgements
We thank Stefano Filipazzi, János Kollár, and Burt Totaro for helpful conversations and comments. Part of this work was carried out while the authors were visiting the Yau Mathematical Sciences Center at Tsinghua University, and we thank Caucher Birkar and Spring Ma for their hospitality.
@@ 2. Preliminaries
We work over an algebraically closed field of characteristic 00. We write $\Sigma^{n}$ for the sum of the coordinate hyperplanes of ${\mathbb{P}}^{n}$. Throughout, a Fano variety will mean a variety $X$ such that $-K_{X}$ is ample and $X$ has klt singularities.
@@ 2.1. Singularities of the MMP
Let $\phi\colon Y\dashrightarrow X$ be a rational map of normal quasi-projective varieties. We say $\phi$ contracts a Weil divisor $E\subset Y$ if the center of $E$ on $X$ is not a divisor. We say $\phi$ extracts a Weil divisor $F\subset X$ if $\phi^{-1}$ contracts $F$. The map $\phi$ is a birational contraction if it is birational and does not extract any divisors.
@@
If $(X,B)$ is an lc pair, a $\mathbb{Q}$-factorial dlt modification of $(X,B)$ is a projective birational morphism $\pi\colon Y\to X$ from a $\mathbb{Q}$-factorial normal variety $Y$ such that ${\rm Ex}(\pi)$ is a divisor, $\pi$ only contracts log canonical places, and the log pullback $(Y,B_{Y})$ is dlt. A $\mathbb{Q}$-factorial dlt modification exists for any lc pair by [32, Theorem 3.1].
@@
Let $X$ be a normal quasi-projective variety.
@@
    We say $X$ is of Fano type if there exists an effective $\mathbb{Q}$-divisor $B$ such that $(X,B)$ is a klt pair and $-(K_{X}+B)$ is big and semiample.
@@
$X$ is Fano type if and only if there exists an effective $\mathbb{Q}$-divisor $B$ on $X$ such that $(X,B)$ is a klt pair and $-(K_{X}+B)$ is ample [44, Lemma-Definition 2.6]. If $X$ is projective, then by [20] an lc pair $(X,B)$ is log CY if and only if $K_{X}+B\equiv 0$.
@@
Two pairs $(X_{1},B_{1})$ and $(X_{2},B_{2})$ are crepant birational equivalent, written $(X_{1},B_{1})\simeq_{\rm cbir}(X_{2},B_{2})$, if there exist proper birational morphisms $f_{i}\colon Y\to X_{i}$ such that the log pullbacks of $B_{i}$ on $Y$ are equal.
@@ 2.2. Toric log Calabi–Yau pairs
Now we define toric and cluster type log Calabi–Yau pairs.
@@
A pair $(X,B)$ is said to be toric if $X$ is a toric variety and $B$ is an effective torus-invariant divisor.
Recall that if $X$ is a normal toric variety and $B$ is the reduced sum of the torus-invariant divisors, then $K_{X}+B\sim 0$ [9, Theorem 8.2.3] and the pair $(X,B)$ is lc [31, Proposition 3.7]. In the case of log CY pairs, the following definitions generalize the notion of being toric.
@@
    We say that $(X,B)$ is of cluster type if there exists a crepant birational map $\phi\colon({\mathbb{P}}^{n},\Sigma^{n})\dashrightarrow(X,B)$ such that $\mathbb{G}_{m}^{n}\cap{\rm Ex}(\phi)$ contains no divisors on $X$.
@@
The reverse implications do not hold; see Example 6.3 and [15, Theorem 3.3 and Proposition 5.1]. If $(X,B)$ is a log CY pair of index one, then $(X,B)$ admits a toric model if and only if the pair $(X,B)$ has birational complexity zero [38, Definition 2.28 and Theorem 1.6]. Since crepant birational equivalences preserve the indices of log CY pairs (see, e.g., [18, Lemma 3.2]), any log rational log CY pair necessarily has index one:
@@
Let $(X,B)$ be a projective toric log CY pair, let $p\colon Y\dashrightarrow X$ be a birational contraction that only contracts non-canonical places of $(X,B)$, and let $(Y,B_{Y})$ be the log pullback of $(X,B)$ to $Y$. Then $(Y,B_{Y})$ is a toric log CY pair, and $Y\dashrightarrow X$ is a toric projective birational morphism.
@@ 2.3. Families of varieties and pairs
Next, we will consider relative versions of Definition 2.7. A projective contraction is a morphism $f\colon X\to T$ of normal quasi-projective varieties such that $f_{*}\cal O_{X}=\cal O_{T}$. A fibration is a projective contraction whose general fiber is positive dimensional.
@@
    A Fano fibration is a projective contraction $X\to T$ such that $-K_{X}$ is ample over $T$.
@@
Let $\Delta$ be a boundary on $X$ such that $(X,\Delta)$ is klt and $-(K_{X}+\Delta)$ is big and semiample over $T$. The log pullback $\Delta_{Y}$ of $\Delta$ is effective as $Y\rightarrow X$ only contracts non-terminal places of $(X,\Delta)$, so $(Y,\Delta_{Y})$ is a klt pair. Since $-(K_{Y}+\Delta_{Y})$ is big and semiample over $T$, the morphism $Y\rightarrow T$ is of Fano type. ∎
The following lemma is well known (see, e.g., [40, Lemma 2.17(3)]).
@@
For a rational map $\pi\colon X\dashrightarrow Z$ of normal varieties, we define the pullback of divisors as follows. Let $\widetilde{X}$ be a normal variety that resolves $\pi$, with projections $p_{1}\colon\widetilde{X}\to X$ and $p_{2}\colon\widetilde{X}\to Z$. Let $D_{Z}$ be a $\mathbb{Q}$-Cartier divisor on $Z$, and assume $mD_{Z}$ is Cartier. Then we define the $\mathbb{Q}$-divisor
@@
[…]
If $D_{X}$ is a prime divisor on $X$, we define its pushforward $\pi_{*}D_{X}$ to be its strict transform on $Z$. We extend $\pi_{*}$ linearly to Weil divisors on $X$. If $\pi$ is a birational contraction, then the equality $\pi_{*}\pi^{*}D=D$ holds as $\mathbb{Q}$-divisors on $Z$. Furthermore, if $\pi$ is a birational contraction, then $\pi_{*}$ induces a homomorphism $\operatorname{Cl}X\to\operatorname{Cl}Z$.
@@ 2.4. Properties of cluster type log Calabi–Yau pairs and dlt modifications
    A formally toric blow-up of $(X^{\rm snc},B^{\rm snc})$ is a blow-up $Y^{\rm snc}\rightarrow X^{\rm snc}$ that is a formally toric morphism over any closed point of $(X^{\rm snc},B^{\rm snc})$ (see [41, Definition 3.42] for more details). We say that a formally toric blow-up is a formally toric dlt blow-up if the log pullback of $(X^{\rm snc},B^{\rm snc})$ to $Y$ is a dlt pair.
@@
Consider the pair $(X,B+\epsilon E)$, which is not necessarily log canonical. By linearity of log discrepancies with respect to the boundary divisor, for $\epsilon>0$ small enough, every non-lc center of $(X,B+\epsilon E)$ is a log canonical center of $(X,B)$. By [32, Theorem 3.1] there exists a $\mathbb{Q}$-factorial dlt modification $\phi\colon(Y,B_{Y}+\epsilon\phi^{-1}_{*}E)\rightarrow(X,B+\epsilon E)$ where $B_{Y}$ is the sum of $\phi^{-1}_{*}B$ and the reduced exceptional divisor of $\phi$ (see [32, Definitions and Notation 1.9]). Then $(Y,B_{Y})$ is dlt, and by the previous considerations, we have $\phi^{*}(K_{X}+B)=K_{Y}+B_{Y}$. Since $(Y,B_{Y}+\epsilon\phi^{-1}_{*}E)$ is log canonical, we know that $\phi^{-1}_{*}E$ contains no log canonical centers of $(Y,B_{Y})$.
Now, assume in addition that $(X,B)$ is dlt. By the previous paragraph, there exists a dlt modification $\phi_{0}\colon(Y_{0},B_{Y_{0}})\rightarrow(X,B)$ such that ${\phi_{0}^{-1}}_{*}E$ contains no log canonical centers of $(Y_{0},B_{Y_{0}})$. We first argue that the dlt modification $(Y_{0},B_{Y_{0}})$ induces a formally toric dlt blow-up $(Y_{0}^{\rm snc},B_{Y_{0}^{\rm snc}})\rightarrow(X^{\rm snc},B^{\rm snc})$. Locally at every closed point, the pair $(X^{\rm snc},B^{\rm snc})$ has local complexity zero [41, Example 3.45, Lemma 3.41, and Definition 3.28]. The morphism $(Y_{0}^{\rm snc},B_{Y^{\rm snc}_{0}})\rightarrow(X^{\rm snc},B^{\rm snc})$ only contracts divisors with coefficient one, so the complexity of this morphism is zero at every closed point of $X^{\rm snc}$ (see [41, Definition 3.15]). Hence, [41, Theorem 1] implies that the dlt modification $(Y_{0}^{\rm snc},B_{Y^{\rm snc}_{0}})\rightarrow(X^{\rm snc},B^{\rm snc})$ is formally toric over the base. That is, $\phi_{0}$ is a formally toric dlt blow-up along some ideal sheaf whose support is contained in ${\rm supp}(B^{\rm snc})$. Now we will construct $Y$ as be a higher model of $Y_{0}^{\rm snc}$ that is obtained from $(X,B)$ by a sequence of formally toric dlt blow-ups along strata.
@@
By Lemma 2.23, every exceptional non-terminal place of $(X,(1-\epsilon)B+\epsilon\Delta))$ is a log canonical place of $(X,B)$. In particular, every exceptional non-terminal place of $(W,\Gamma_{W})$ is a log canonical place of $(W,B_{W})$, and the log canonical centers of $(W,B_{W})$ are precisely its proper strata. By construction, the components of $\Gamma_{W}$ that are not contained in $B_{W}$ are precisely the components of $\phi^{-1}_{*}D$. So, on a neighborhood of the generic point of every proper stratum of $(W,B_{W})$, the support of $\Gamma_{W}$ is contained in the support of $B_{W}$. Since $(W,B_{W})$ is log smooth along the generic point of each stratum, we conclude that $(W,\Gamma_{W})$ is log smooth around every non-terminal center. Therefore, we may obtain a terminal model of $(W,\Gamma_{W})$ by successively blowing up strata of $(W,B_{W})$ (see [34, Lemma 2.29 and Proposition 2.36(2)]). Hence, the composition $(Y,B_{Y})\rightarrow(W,B_{W})\rightarrow(Z,B_{Z})$ is a $\mathbb{Q}$-factorial dlt modification for which the log pullback of $(X,(1-\epsilon)B+\epsilon\Delta)$ is terminal. ∎
@@
[…]
@@
Since $(Y,B)$ and $(Y_{t},B_{t})$ are dlt, the lc centers of $(Y_{t},B_{t})$ are irreducible components of restrictions of lc centers of $(Y,B)$. So, by assumption, $E$ contains no lc centers of $(Y_{t},B_{t})$. Then inversion of adjunction [24, Theorem 1] and the assumption that $Y_{t}\not\subset E$ imply that $E$ contains no log canonical centers of $(Y,B+f^{*}H_{1}+\dots+f^{*}H_{d})$ near $t$. Thus, for sufficiently small $\delta>0$, the pair $(Y,B+f^{*}H_{1}+\dots+f^{*}H_{d}+\delta E)$ is log canonical over a neighborhood of $t$.
@@
Since the log fiber $(W_{t},B_{t})$ is the pair obtained by adjunction of (2.7) to $W_{t}$, we conclude by [24, Theorem 1] that its normalization is lc. ∎

### v2

@@ 1. Introduction
Let $\mathcal{X}\to T$ be a family of projective varieties over $\mathbb{C}$. The subset
[…]
@@
In fact, the Fano assumption is not necessary: being toric is a constructible condition in any family of complex varieties (Theorem 3.4). We note that being toric is neither an open nor a closed condition (Examples 6.1 and 6.2).
Being toric places strong restrictions on the geometry of a variety. For instance, smooth toric Fano varieties do not deform [2] (see also [11, Theorem 1.4]), so Theorem 1.1 is automatic in the smooth case. Thus, we consider an intermediate class: cluster type varieties (see Definition 2.11). This is a class of rational varieties generalizing toric varieties, and was introduced by Enwright, Figueroa, and the second author in [15]. For example, smooth toric varieties are covered by tori, up to a closed codimension two subset (see Proposition 2.9), whereas smooth cluster type Fano varieties contain an open dense torus (see [15, Theorem 1.2 and Theorem 1.3(4)]). We prove the following result about cluster type varieties in families of Fano varieties:
@@
By parametrizing, we mean here that every $n$-dimensional smooth cluster type Fano appears as the closed fiber of some $f_{i}$, and, moreover, every closed fiber of each $f_{i}$ is in this class of varieties. This notion is stronger than boundedness, but weaker than moduli, since the same variety could appear as a fiber multiple times. Furthermore, from the proof of Theorem 1.2, it follows that all fibers of each $f_{i}$ in Corollary 1.3 admit the same combinatorial structure, in the sense that a certain birational transformation to a toric pair can be performed in the whole family (see Lemma 2.30).
On the other hand, the analogous statement for rational varieties is false:
Theorem 1.4.
Let $n\geq 4$ be an integer. Then, there do not exist finitely many smooth projective morphisms parametrizing $n$-dimensional smooth rational Fano varieties.
In the toric case, the analogous statement to Corollary 1.3 was already known: smooth toric Fano varieties correspond to smooth reflexive polytopes (see, e.g., [9, Theorem 8.3.4]), so in each dimension $n$ there are only finitely many up to isomorphism. Note that smooth cluster type Fanos do deform, so unlike in the toric case, the parametrizing varieties $T_{i}$ in Corollary 1.3 are in general positive-dimensional. For example, every smooth del Pezzo surface of degree $\geq 2$ is of cluster type by [1, Theorem 2.1] and [22, Lemma 1.13].
@@ Acknowledgements
We are grateful to János Kollár and Burt Totaro for suggesting Theorem 3.4, and to Brendan Hassett for suggesting Theorem 1.4. We also thank Louis Esser and Stefano Filipazzi for helpful conversations and comments, and we thank the anonymous referee for their careful reading and helpful comments. Part of this work was carried out while the authors were visiting the Yau Mathematical Sciences Center at Tsinghua University, and we thank Caucher Birkar and Spring Ma for their hospitality.
@@ 2. Preliminaries
We work over an algebraically closed field of characteristic 0. We write $\Sigma^{n}$ for the sum of the coordinate hyperplanes of ${\mathbb{P}}^{n}$. Throughout, a Fano variety will mean a projective variety $X$ such that $K_{X}$ is ample and $X$ has klt singularities.
@@ 2.1. Singularities of the MMP
Let $\phi\colon Y\dashrightarrow X$ be a rational map of normal projective varieties. We say $\phi$ contracts a Weil divisor $E\subset Y$ if the center of $E$ on $X$ is not a divisor. We say $\phi$ extracts a Weil divisor $F\subset X$ if $\phi^{-1}$ contracts $F$. The map $\phi$ is a birational contraction if it is birational and does not extract any divisors.
@@
If $(X,B)$ is an lc pair, a $\mathbb{Q}$-factorial dlt modification of $(X,B)$ is a projective birational morphism $\pi\colon Y\to X$ from a $\mathbb{Q}$-factorial normal variety $Y$ such that ${\rm Ex}(\pi)$ is a divisor, $\pi$ only contracts log canonical places, and the log pullback $(Y,B_{Y})$ is dlt. A $\mathbb{Q}$-factorial dlt modification exists for any lc pair by [34, Theorem 3.1].
@@
Let $X$ be a normal projective variety.
@@
    We say $X$ is of Fano type if there exists an effective $\mathbb{Q}$-divisor $B$ such that $(X,B)$ is a klt pair and $(K_{X}+B)$ is big and semiample.
@@
$X$ is Fano type if and only if there exists an effective $\mathbb{Q}$-divisor $B$ on $X$ such that $(X,B)$ is a klt pair and $(K_{X}+B)$ is ample [46, Lemma-Definition 2.6]. By [20] an lc pair $(X,B)$ is log CY if and only if $K_{X}+B\equiv 0$.
@@
Two pairs $(X_{1},B_{1})$ and $(X_{2},B_{2})$ are crepant birational equivalent, written $(X_{1},B_{1})\simeq_{\rm cbir}(X_{2},B_{2})$, if there exist projective birational morphisms $f_{i}\colon Y\to X_{i}$ such that the log pullbacks of $B_{i}$ on $Y$ are equal.
@@ 2.2. Toric log Calabi–Yau pairs
We will next define toric and cluster type log Calabi–Yau pairs. Before this, we first prove a proposition regarding a statement mentioned in the introduction:
Proposition 2.9.
Let $T$ be a normal quasi-projective toric variety of dimension $n$. Then, there exists a torus-invariant closed subset $Z\subset T$, of codimension at least two, such that $T\setminus Z$ is covered by copies of $\mathbb{G}_{m}^{n}$.
Proof.
Let $\Sigma$ be the fan in ${\mathbb{Q}}^{n}$ defining $T$, i.e., $T\simeq X(\Sigma)$. By [9, Theorem 3.2.6(a)], $T$ has finitely many torus-invariant closed subsets of codimension at least two, so their union $Z$ is a closed subset of $T$ of codimension at least two. We argue that $T\setminus Z$ is covered by copies of $\mathbb{G}_{m}^{n}$. Let $\Sigma(1)$ be the set of rays of $\Sigma$, and for $\rho\in\Sigma(1)$, let $D_{\rho}$ be the associated torus-invariant prime divisor. Now, fix an order $\rho_{1},\dots,\rho_{k}$ of the elements of $\Sigma(1)$. Then, for each $j\in\{1,\dots,k\}$, we have
$T\setminus\bigcup_{i\neq j}D_{\rho_{i}}=\bigcup_{\tau\preceq\rho_{i}}O(\tau)\simeq\mathbb{G}_{m}^{n-1}\times\mathbb{A}^{1}$
by [9, Theorem 3.2.6(c)] because $\rho_{i}$ is a ray. This implies that $T\setminus Z$ is covered by copies of $\mathbb{G}_{m}^{n}$. ∎
@@
A pair $(X,B)$ is said to be toric if $X$ is a projective toric variety and $B$ is a torus-invariant divisor.
Recall that if $X$ is a normal toric variety and $B$ is the reduced sum of the torus-invariant divisors, then $K_{X}+B\sim 0$ [9, Theorem 8.2.3] and the pair $(X,B)$ is lc [32, Proposition 3.7]. In the case of log CY pairs, the following definitions generalize the notion of being toric.
@@
    We say that $(X,B)$ is of cluster type if there exists a crepant birational map $\phi\colon({\mathbb{P}}^{n},\Sigma^{n})\dashrightarrow(X,B)$ such that $\mathbb{G}_{m}^{n}\cap{\rm Ex}(\phi)$ contains no divisors on $\mathbb{P}^{n}$.
@@
The reverse implications do not hold; see Example 6.3 and [15, Theorem 3.3 and Proposition 5.1]. If $(X,B)$ is a log CY pair of index one, then $(X,B)$ admits a toric model if and only if the pair $(X,B)$ has birational complexity zero [40, Definition 2.28 and Theorem 1.6]. Since crepant birational equivalences preserve the indices of log CY pairs (see, e.g., [18, Lemma 3.2]), any log rational log CY pair necessarily has index one:
@@
Let $(X,B)$ be a projective toric log CY pair, let $p\colon Y\dashrightarrow X$ be a birational contraction that only contracts non-canonical places of $(X,B)$, and let $(Y,B_{Y})$ be the log pullback of $(X,B)$ to $Y$. Then $(Y,B_{Y})$ is a toric log CY pair, and $Y\dashrightarrow X$ is a toric projective birational contraction.
@@ 2.3. Families of varieties and pairs
Next, we will consider relative versions of Definition 2.7. A projective contraction is a projective morphism $f\colon X\to T$ of normal quasi-projective varieties such that $f_{*}\mathcal{O}_{X}=\mathcal{O}_{T}$. A fibration is a projective contraction whose general fiber is positive dimensional.
@@
    A Fano fibration is a projective contraction $X\to T$ such that $K_{X}$ is ample over $T$.
@@
Let $\Delta$ be a boundary on $X$ such that $(X,\Delta)$ is klt and $(K_{X}+\Delta)$ is big and semiample over $T$. The log pullback $\Delta_{Y}$ of $\Delta$ is effective as $Y\rightarrow X$ only contracts non-terminal places of $(X,\Delta)$, so $(Y,\Delta_{Y})$ is a klt pair. Since $(K_{Y}+\Delta_{Y})$ is big and semiample over $T$, the morphism $Y\rightarrow T$ is of Fano type. ∎
The following lemma is well known (see, e.g., [42, Lemma 2.17(3)]).
@@
For a rational map $\pi\colon X\dashrightarrow Z$ of normal projective varieties, we define the pullback of divisors as follows. Let $\widetilde{X}$ be a normal variety that resolves $\pi$, with projections $p_{1}\colon\widetilde{X}\to X$ and $p_{2}\colon\widetilde{X}\to Z$. Let $D_{Z}$ be a $\mathbb{Q}$-Cartier divisor on $Z$, and assume $mD_{Z}$ is Cartier. Then we define the $\mathbb{Q}$-divisor
@@
$H^{0}(Z,\mathcal{O}_{Z}(mD_{Z}))\xrightarrow{\cong}H^{0}(Z,{p_{2}}_{*}p_{2}^{*}\mathcal{O}_{Z}(mD_{Z}))=H^{0}(\widetilde{X},p_{2}^{*}\mathcal{O}_{Z}(mD_{Z})).$
If $D_{X}$ is a prime divisor on $X$, we define its pushforward $\pi_{*}D_{X}$ to be its strict transform on $Z$. We extend $\pi_{*}$ linearly to Weil divisors on $X$. If $\pi$ is a birational contraction, then the equality $\pi_{*}\pi^{*}D_{Z}=D_{Z}$ holds as $\mathbb{Q}$-divisors on $Z$. Furthermore, if $\pi$ is a birational contraction, then $\pi_{*}$ induces a homomorphism $\operatorname{Cl}X\to\operatorname{Cl}Z$.
@@ 2.4. Properties of cluster type log Calabi–Yau pairs and dlt modifications
    A formally toric blow-up of $(X^{\rm snc},B^{\rm snc})$ is a blow-up $Y^{\rm snc}\rightarrow X^{\rm snc}$ that is a formally toric morphism over any closed point of $(X^{\rm snc},B^{\rm snc})$ (see [43, Definition 3.42] for more details). We say that a formally toric blow-up is a formally toric dlt blow-up if the log pullback of $(X^{\rm snc},B^{\rm snc})$ to $Y$ is a dlt pair.
@@
Consider the pair $(X,B+\epsilon E)$, which is not necessarily log canonical. By linearity of log discrepancies with respect to the boundary divisor, for $\epsilon>0$ small enough, every non-lc center of $(X,B+\epsilon E)$ is a log canonical center of $(X,B)$. By [34, Theorem 3.1] there exists a $\mathbb{Q}$-factorial dlt modification $\phi\colon(Y,B_{Y}+\epsilon\phi^{-1}_{*}E)\rightarrow(X,B+\epsilon E)$ where $B_{Y}$ is the sum of $\phi^{-1}_{*}B$ and the reduced exceptional divisor of $\phi$ (see [34, Definitions and Notation 1.9]). Then $(Y,B_{Y})$ is dlt, and by the previous considerations, we have $\phi^{*}(K_{X}+B)=K_{Y}+B_{Y}$. Since $(Y,B_{Y}+\epsilon\phi^{-1}_{*}E)$ is log canonical, we know that $\phi^{-1}_{*}E$ contains no log canonical centers of $(Y,B_{Y})$.
Now, assume in addition that $(X,B)$ is dlt. By the previous paragraph, there exists a dlt modification $\phi_{0}\colon(Y_{0},B_{Y_{0}})\rightarrow(X,B)$ such that ${\phi_{0}^{-1}}_{*}E$ contains no log canonical centers of $(Y_{0},B_{Y_{0}})$. We first argue that the dlt modification $(Y_{0},B_{Y_{0}})$ induces a formally toric dlt blow-up $(Y_{0}^{\rm snc},B_{Y_{0}^{\rm snc}})\rightarrow(X^{\rm snc},B^{\rm snc})$. Locally at every closed point, the pair $(X^{\rm snc},B^{\rm snc})$ has local complexity zero [43, Example 3.45, Lemma 3.41, and Definition 3.28]. The morphism $(Y_{0}^{\rm snc},B_{Y^{\rm snc}_{0}})\rightarrow(X^{\rm snc},B^{\rm snc})$ only contracts divisors with coefficient one, so the complexity of this morphism is zero at every closed point of $X^{\rm snc}$ (see [43, Definition 3.15]). Hence, [43, Theorem 1] implies that the dlt modification $(Y_{0}^{\rm snc},B_{Y^{\rm snc}_{0}})\rightarrow(X^{\rm snc},B^{\rm snc})$ is formally toric over the base. That is, $\phi_{0}$ is a formally toric dlt blow-up along some ideal sheaf whose support is contained in ${\rm supp}(B^{\rm snc})$. Now we will construct $Y$ as a higher model of $Y_{0}^{\rm snc}$ that is obtained from $(X,B)$ by a sequence of formally toric dlt blow-ups along strata.
@@
By Lemma 2.24, every exceptional non-terminal place of $(X,(1-\epsilon)B+\epsilon\Delta))$ is a log canonical place of $(X,B)$. In particular, every exceptional non-terminal place of $(W,\Gamma_{W})$ is a log canonical place of $(W,B_{W})$, and the log canonical centers of $(W,B_{W})$ are precisely its proper strata. By construction, the components of $\Gamma_{W}$ that are not contained in $B_{W}$ are precisely the components of $\phi^{-1}_{*}D$. So, on a neighborhood of the generic point of every proper stratum of $(W,B_{W})$, the support of $\Gamma_{W}$ is contained in the support of $B_{W}$. Since $(W,B_{W})$ is log smooth along the generic point of each stratum, we conclude that $(W,\Gamma_{W})$ is log smooth around every non-terminal center. Therefore, we may obtain a terminal model of $(W,\Gamma_{W})$ by successively blowing up strata of $(W,B_{W})$ (see [36, Lemma 2.29 and Proposition 2.36(2)]). Hence, the composition $(Y,B_{Y})\rightarrow(W,B_{W})\rightarrow(Z,B_{Z})$ is a $\mathbb{Q}$-factorial dlt modification for which the log pullback of $(X,(1-\epsilon)B+\epsilon\Delta)$ is terminal. ∎
@@
${(Y_{U},\Gamma_{U}+f^{*}H_{1}+\dots+f^{*}H_{d}+\delta E)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$$\scriptstyle{f}$${(W,\Gamma_{W}+{f^{\prime}}^{*}H_{1}+\dots+{f^{\prime}}^{*}H_{d}+\delta E_{W})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}}$${T}$
@@
Since $(Y,B)$ and $(Y_{t},B_{t})$ are dlt, the lc centers of $(Y_{t},B_{t})$ are irreducible components of restrictions of lc centers of $(Y,B)$. So, by assumption, $E$ contains no lc centers of $(Y_{t},B_{t})$. Then inversion of adjunction [25, Theorem 1] and the assumption that $Y_{t}\not\subset E$ imply that $E$ contains no log canonical centers of $(Y,B+f^{*}H_{1}+\dots+f^{*}H_{d})$ near $t$. Thus, for sufficiently small $\delta>0$, the pair $(Y,B+f^{*}H_{1}+\dots+f^{*}H_{d}+\delta E)$ is log canonical over a neighborhood of $t$.
@@
Since the log fiber $(W_{t},B_{t})$ is the pair obtained by adjunction of (2.7) to $W_{t}$, we conclude by [25, Theorem 1] that its normalization is lc. ∎